# Teleportation=Translation: Continuous recovery of black hole information

### v2

@@ 5 Conclusion
We have presented a rigorous bridge between the discrete, algebraic teleportation protocol of van den Heijden and Verlinde vdHV and the continuous, geometric picture provided by modular flow. In Type III settings, the key difficulty is to construct a path that is both mathematically controlled and physically consistent. Our construction achieves this and leads to a strongly continuous unitary path $\tilde{U}(s)$ with a well-defined self-adjoint generator $\tilde{G}$, obeying the relation
@@
The unitary $\tilde{U}(s)\approx e^{-is\tilde{G}}$ implements a literal motion of quantum information. At $s=0$, the information is dynamically inaccessible, residing in the relative commutant of the inclusion $\mathcal{N}\subset\mathcal{M}$ ($\mathcal{A}=\mathcal{N}^{\prime}\cap\mathcal{M}$, corresponding to the island region just behind the horizon). As $s$ increases, the operator smoothly transports this information outwards into the accessible radiation sector. The endpoint $\Gamma=\tilde{U}(1)$ is the canonical shift, and the factor of $2$ arises because $\Gamma$ is generated by two modular conjugations, effectively doubling the modular distance. Thus, the act of recovering information from the island via teleportation is algebraically identical to a geometric spacetime translation generated by $P$, making Teleportation=Translation precise.
This viewpoint yields a concrete resolution of the black hole information paradox. Global evolution remains unitary. $\tilde{U}(s)$ transports, rather than destroys, interior information to the exterior algebra, so no information is lost during evaporation. No-cloning is preserved by algebraic complementarity: the same information is described in different algebras at different $s$, never duplicated at once. Apparent thermality is a property of Type III subalgebras (absence of a tracial state), not of the global dynamics; lifting to the semifinite envelope clarifies that entropy and thermality are governed by modular, not tracial, structure. Dynamic idempotency gives a nested flow of algebras, providing a controlled, continuous version of the Page/Hayden–Preskill decoding protocol. In holographic and HSMI contexts, where $P=K_{\mathcal{M}}-K_{\mathcal{N}}$ generates horizon translation, the equality $\tilde{G}=2P$ identifies the recovery unitary with a geometric translation in the emergent spacetime.
@@
This condensed, physics-forward account emphasizes that continuous, idempotent, modular-covariant teleportation is spacetime translation, providing a unitary and calculable information-transfer channel through which black holes evaporate without violating quantum mechanics or algebraic locality.
It is important to emphasize that these results are established within an algebraic QFT framework on a fixed background spacetime. While we demonstrate that information flow remains unitary and consistent with geometric translation in this setting, this work provides the kinematic skeleton of information recovery; the full quantum gravitational dynamics, including the back-reaction of the evaporation process on the geometry, remains an open frontier.
Our work points to several possible directions for future research. First, the connection between $\tilde{G}=2P$ and traversable wormholes warrants deeper investigation. Since $P$ generates translations across the horizon, our continuous protocol may provide an algebraic description of the Gao-Jafferis-Wall protocol Gao2017 , potentially offering a non-perturbative definition of the wormhole opening mechanism via modular flow. Second, explicit verification of the correlation function equality $F^{\prime}(0)\propto-i\langle[P,\mathcal{O}]\rangle$ in holographic models, such as double-scaled Sachdev-Ye-Kitaev (DSSYK) Berkooz:2018jqr or Jackiw-Teitelboim (JT) gravity Saad:2019lba , would provide a direct bridge between our algebraic results and semiclassical gravity calculations.
Finally, extending this framework to include gravitational back-reaction is a critical next step. Recent developments suggest that including $1/N$ corrections or observer energy naturally leads to a transition from Type III to Type II algebras via the crossed product construction Witten22 ; Chandrasekaran2022cip ; LeutheusserLiu_ET . This transition mathematically implements the gravitational constraints, thereby capturing the leading-order back-reaction of quantum matter on the geometry. Since our protocol $\tilde{U}(s)$ is explicitly constructed within this crossed-product envelope, it provides a natural starting point for incorporating back-reaction. Formulating the interplay between the information transport $\tilde{U}(s)$ and the semi-classical Einstein equations in this Type II setting could pave the way for a fully dynamic theory of emergent spacetime geometry from quantum entanglement.
@@ Appendix A Notation table
$\mathcal{M},\mathcal{N}$ von Neumann algebras ($\mathcal{N}\subset\mathcal{M}$) for the two regions/systems studied
$\tilde{\mathcal{M}},\tilde{\mathcal{N}}$ semifinite crossed-product envelopes (Haagerup–Takesaki) algebras
$\mathcal{M}^{+},\mathcal{N}^{+}$ positive cones of $\mathcal{M},\mathcal{N}$, respectively
@@
$J_{\mathcal{X}},\Delta_{\mathcal{X}},K_{\mathcal{X}}$ modular conjugation, modular operator, modular Hamiltonian for algebra $\mathcal{X}$
$U_{\Gamma}$ discrete canonical shift (van der Heijden–Verlinde shift)
@@
$P$ shorthand for the modular Hamiltonian difference $K_{\tilde{\mathcal{M}}}-K_{\tilde{\mathcal{N}}}$; $2P$ is the claimed generator
$\phi_{0},\phi_{1}$ faithful normal weights on $\tilde{\mathcal{M}}$ associated with the algebras (see Sec. 4.2)
@@
Appendix B Type I Model: Path Failure and Unitary Interpolation Goal
@@
such that $U(0)=\mathbf{1}$ (the identity) and $U(1)\propto U_{\Gamma}$ (coinciding with the swap up to a global phase). This illustrates the physical goal of the construction: finding a continuous time evolution $U(s)$ that connects the initial state to the final shift, providing a valid unitary path even though the corresponding conditional expectation path is mathematically problematic in the Type III limit.
@@ Appendix C Haagerup–Kosaki Lift: Existence and Mechanism of $\tilde{E}$
The most critical technical challenge in defining the idempotent path $\tilde{E}_{s}$ in the Type III setting is the construction of a faithful normal conditional expectation. In general, Type III inclusions do not admit a conditional expectation $\mathcal{M}\to\mathcal{N}$. The theory of Haagerup and Takesaki provides the necessary mechanism by lifting the problem from a non-tracial Type III algebra (equipped with an OVW) to a tracial semifinite algebra where a genuine conditional expectation $\tilde{E}$ exists. This lift effectively regularizes the unbounded OVW $\mathcal{E}$ via the crossed product construction.
@@
Let $\mathcal{N}\subset\mathcal{M}$ be an inclusion of von Neumann algebras. While a conditional expectation may not exist, Haagerup’s theorem guarantees the existence of a faithful normal OVW $\mathcal{E}:\mathcal{M}^{+}\to\mathcal{N}^{+}$.
@@
(This condition is equivalent to requiring $\omega=\psi\circ\mathcal{E}$ for some weight $\psi$ on $\mathcal{N}$).
@@
    The subalgebra inclusion lifts to $\tilde{\mathcal{N}}=\mathcal{N}\rtimes_{\sigma^{\omega}|_{\mathcal{N}}}\mathbb{R}\subset\tilde{\mathcal{M}}$.
@@
    Takesaki Duality and Trace: Takesaki duality establishes that $\tilde{\mathcal{M}}$ is a semifinite algebra equipped with a canonical trace $\tau$. The scaling property (38) is characteristic of the Type II crossed product derived from a Type III algebra, allowing us to treat modular dynamics as trace-scaling automorphisms. Consequently, the existence of this trace allows for the existence of conditional expectations that are not available in the Type III setting.
@@
    Construction of $\tilde{E}$: The existence of the trace-preserving conditional expectation $\tilde{E}$ follows from the compatibility of the dual weights on $\tilde{\mathcal{M}}$ and $\tilde{\mathcal{N}}$. Specifically, because the original weights were chosen to be compatible ($\omega=\psi\circ\mathcal{E}$), their dual weights on the crossed products relate via a conditional expectation. This $\tilde{E}$ effectively "renormalizes" the original OVW $\mathcal{E}$ into a norm-one projection on the semifinite algebra.
@@
    Equivariance: The construction explicitly preserves the structure of the dual action $\theta_{t}$, ensuring $\tilde{E}\circ\theta_{t}=\theta_{t}\circ\tilde{E}$. This property is crucial for proving that the analytic core is preserved under the operation of $\tilde{E}$ (as discussed in Section 3.4).

### v3

@@ 5 Conclusion
We have presented a framework connecting the discrete, algebraic teleportation protocol of vdH-V vdHV with the continuous, geometric picture provided by modular flow. In Type III settings, a key challenge is to construct a path that is both mathematically controlled and physically consistent. Our construction addresses this by defining a strongly continuous unitary path $\tilde{U}(s)$ with a well-defined self-adjoint generator $\tilde{G}$, satisfying the relation
@@
The unitary $\tilde{U}(s)\approx e^{-is\tilde{G}}$ describes an effective transport of quantum information. At $s=0$, the information resides in the relative commutant $\tilde{\mathcal{A}}=\tilde{\mathcal{N}}^{\prime}\cap\tilde{\mathcal{M}}$ (corresponding to the island region just behind the horizon) and is dynamically inaccessible. As $s$ increases, the operator smoothly transports this information into the accessible radiation sector $\tilde{\mathcal{M}}^{\prime}\cap\tilde{\mathcal{M}}_{1}(s)$, where $\tilde{\mathcal{M}}_{1}(s):=\langle\tilde{\mathcal{M}},e_{\tilde{\mathcal{N}}(s)}\rangle$. The endpoint corresponds to the canonical shift unitary $U_{\tilde{\Gamma}}$, where the factor of $2$ arises because $U_{\tilde{\Gamma}}$ is generated by two modular conjugations, effectively doubling the modular distance. In this view, the recovery of information via teleportation becomes algebraically analogous to a geometric spacetime translation generated by $P$, giving a precise form to the correspondence Teleportation=Translation.
This perspective offers a consistent framework for approaching the black hole information paradox. Within this picture, global evolution is unitary: the operator $\tilde{U}(s)$ transports interior information to the exterior algebra rather than destroying it, suggesting that no information is lost during the evaporation process. No-cloning is naturally preserved by algebraic complementarity, as the information is encoded in distinct algebras across different values of $s$ without duplication. The apparent thermality can be interpreted as a consequence of the Type III algebraic structure—specifically the absence of a tracial state—rather than an intrinsic feature of the global dynamics. Furthermore, lifting the description to the semifinite envelope clarifies that entropy and thermality are governed by modular, as opposed to tracial, structures.
The property of dynamic idempotency induces a nested flow of algebras, offering a controlled, continuous realization of the Page/Hayden–Preskill decoding protocol. This framework shares key conceptual features with holographic quantum error correction and entanglement wedge reconstruction. In particular, the smooth transport of information from the initially inaccessible relative commutant to the radiation sector is analogous to the recoverability conditions in error-correcting codes Almheiri_2015 . Similarly, the generator $\tilde{G}=2P$ provides an algebraic description of accessing the island interior, paralleling the geometric reconstruction of the entanglement wedge Penington19 .
@@
In summary, this formal development indicates that continuous, idempotent, and modular-covariant teleportation is algebraically equivalent to a spacetime translation. This identification provides a unitary and calculable channel for information transfer, supporting the view that black holes may evaporate in accordance with the principles of quantum mechanics and algebraic locality.
It is worth noting that these results are derived within an algebraic QFT framework on a fixed background spacetime. While we demonstrate that information flow remains unitary and consistent with geometric translation in this setting, this work primarily provides a kinematic framework for information recovery; the full quantum gravitational dynamics, including the back-reaction of the evaporation process on the geometry, remains an open and important frontier.
Our work suggests several directions for future research. First, the connection between $\tilde{G}=2P$ and traversable wormholes warrants deeper investigation. Since $P$ generates translations across the horizon, our continuous protocol may provide an algebraic description of the GJW protocol Gao2017 , potentially offering a non-perturbative perspective on the wormhole opening mechanism via modular flow. Second, verifying the correlation function equality $F^{\prime}(0)\propto-i\langle[P,\mathcal{O}]\rangle$ in holographic models, such as double-scaled Sachdev-Ye-Kitaev (DSSYK) Berkooz:2018jqr or JT gravity, would help bridge our algebraic results with semiclassical gravity calculations.
Finally, extending this framework to incorporate gravitational back-reaction is a critical next step. Since our protocol $\tilde{U}(s)$ is explicitly constructed within the crossed-product envelope—a structure known to capture $1/N$ corrections and observer energy constraints Witten22 ; Chandrasekaran2022cip ; LeutheusserLiu_ET —it offers a natural starting point for such an extension. Formulating the interplay between the information transport $\tilde{U}(s)$ and the semi-classical Einstein equations in this Type II setting could pave the way for a dynamic theory of emergent spacetime geometry from quantum entanglement.
@@ Appendix A Notation table
$\mathcal{M},\ \mathcal{N}$ von Neumann algebras ($\mathcal{N}\subset\mathcal{M}$)
$\tilde{\mathcal{M}},\ \tilde{\mathcal{N}}$ semifinite crossed-product algebras (envelopes)
$\mathcal{X}^{+}$, $\widehat{\mathcal{X}}^{+}$ positive and extended positive cones of $\mathcal{X}$
[…]
@@
$U_{\Gamma}$, $U_{\tilde{\Gamma}}$ discrete and lifted canonical shift unitaries
$\quad\tilde{\mathcal{A}}$ relative commutant $\tilde{\mathcal{N}}^{\prime}\cap\tilde{\mathcal{M}}$ (island algebra)
$\tilde{\mathcal{M}}_{1}(s)$ basic extension algebra $\langle\tilde{\mathcal{M}},e_{\tilde{\mathcal{N}}(s)}\rangle$
$\tilde{\mathcal{M}}^{\prime}\cap\tilde{\mathcal{M}}_{1}(s)$ accessible radiation sector (basic extension’s commutant)
@@
$h=d\tilde{\phi}_{1}/d\tilde{\phi}_{0}$ non-commutative Radon–Nikodym derivative
$K_{\mathcal{X}}$, $P$ modular Hamiltonian of $\mathcal{X}$ and modular momentum (difference of Hamiltonians)
$\quad\sigma_{t}^{\phi}$ modular automorphism group of weight $\phi$
$J_{\mathcal{X}}$, $\Delta_{\mathcal{X}}$ modular conjugation and modular operator
$\quad\tau$ canonical trace on the crossed product $\tilde{\mathcal{M}}$
$\tilde{\phi}_{0}$, $\tilde{\phi}_{1}$ faithful normal weights on $\tilde{\mathcal{M}}$
@@
Appendix B Finite-Dimensional Model: Idempotency Failure and Unitary Interpolation
@@
such that $U(0)=\mathbf{1}$. At $s=1$, we find:
$U(1)=e^{-i\frac{\pi}{2}\sigma_{x}}=\cos(\frac{\pi}{2})\mathbf{1}-i\sin(\frac{\pi}{2})\sigma_{x}=-i\sigma_{x}\propto U_{\Gamma}$ (40)
Thus, the endpoint coincides with the swap up to a global phase factor of i-i. This illustrates the physical goal of the construction: finding a continuous evolution $U(s)$ that connects the initial state to the final shift. This unitary path serves as a consistent alternative to the naive CP-map interpolation, which was shown to be mathematically problematic in Example B.1.
@@ Appendix C Haagerup–Kosaki Lift: Existence and Mechanism of $\tilde{E}$
The most critical technical challenge in defining the idempotent path $E_{s}$ in the Type III setting is the construction of a faithful normal conditional expectation. In general, Type III inclusions do not admit a conditional expectation $\mathcal{M}\to\mathcal{N}$. The theory of Haagerup and Takesaki provides the necessary mechanism by lifting the problem from a non-tracial Type III algebra (equipped with an operator-valued weight, OVW) to a tracial semifinite algebra where a genuine conditional expectation $\tilde{E}$ exists. This lift effectively regularizes the unbounded OVW $\mathcal{E}$ via the crossed product construction.
@@
Let $\mathcal{N}\subset\mathcal{M}$ be an inclusion of Type III von Neumann algebras. While a conditional expectation may not exist, Haagerup’s theorem guarantees the existence of a faithful normal OVW $\mathcal{E}:\mathcal{M}^{+}\to\widehat{\mathcal{N}}^{+}$.
@@
(This condition implies that $\omega$ can be written as $\omega=\psi\circ\mathcal{E}$ for some faithful normal weight $\psi$ on $\mathcal{N}$).
@@
    Furthermore, the connection to the base algebra is maintained by the canonical OVW $\underline{\mathcal{E}}:\tilde{\mathcal{M}}^{+}\to\widehat$$\mathcal{M}$$^{+}$ (often denoted as the dual weight construction), formally given by:
[…]
@@
    The subalgebra inclusion lifts to $\tilde{\mathcal{N}}=\mathcal{N}\rtimes_{\sigma^{\omega}|_{\mathcal{N}}}\mathbb{R}\subset\tilde{\mathcal{M}}$. Crucially, this lifted subalgebra remains invariant under the dual action:
    $\theta_{t}(\tilde{\mathcal{N}})=\tilde{\mathcal{N}}$. (44)
@@
    Semifinite Structure and Trace: The theory of crossed products guarantees that $\tilde{\mathcal{M}}$ is a semifinite algebra equipped with a canonical trace $\tau$. The scaling property (42) is a defining characteristic of the Type II crossed product derived from a Type III algebra, allowing us to treat modular dynamics as trace-scaling automorphisms. Consequently, the existence of this trace allows for the construction of conditional expectations that are not available in the Type III setting. ă
@@
    Construction of $\tilde{E}$: The existence of the trace-preserving conditional expectation $\tilde{E}$ follows from the compatibility of the dual weights on $\tilde{\mathcal{M}}$ and $\tilde{\mathcal{N}}$. Specifically, because the original weights were chosen to be compatible ($\omega=\psi\circ\mathcal{E}$), their dual traces on the crossed products coincide, implying the existence of a conditional expectation. This $\tilde{E}$ effectively renormalizes the original OVW $\mathcal{E}$ into a norm-one projection on the semifinite algebra. ă
@@
    Equivariance: The construction explicitly preserves the structure of the dual action $\theta_{t}$, ensuring $\tilde{E}\circ\theta_{t}=\theta_{t}\circ\tilde{E}$. This property is crucial for proving that the analytic domains are preserved under the operation of $\tilde{E}$.